OpenStudy (anonymous):

Derivative of (sqrt(3)/4) s^2? I am using the product rule... f(x) g'(x) + g(x) f'(x).... but I'm stuck with (sqrt(3)/4). What's the derivative of (sqrt(3)/4)?

OpenStudy (usukidoll):

0

OpenStudy (usukidoll):

there's no variable.... OH! now I see it sorry

OpenStudy (anonymous):

thats okay

OpenStudy (usukidoll):

HUH! why product rule? chain rule is more like it... soorryy I need to getmy homework done. ACK!

OpenStudy (usukidoll):

no wrong I SORRY ^^

OpenStudy (anonymous):

Chain rule?

OpenStudy (anonymous):

i don't think it's supposed to chain rule... don't you take the derivative of each?

OpenStudy (usukidoll):

(sqrt(3)/4) s^2 leave sqrt(3/4) alone deal with s^2 that's 2s + leave s^2 alone deal with sqrt (3/4) that's just a number without a variable so it's 0. multiply this part and that's 0. 2s sqrt(3/4)

OpenStudy (usukidoll):

number a.k.a constant... no derivatives..they go to 0

OpenStudy (anonymous):

it's not sqrt(3/4). It's sqrt(3) divided by 4

OpenStudy (usukidoll):

kind of hard to see when draw is broken on os eh

OpenStudy (anonymous):

i wish i can use the equation tool... it's not working again

OpenStudy (superdavesuper):

d/dx(k*f(x)) = k*d/dx(f(x)) where k is a constant so just move the sqrt part out from ur derivative...

OpenStudy (anonymous):

what do you mean? so sqrt(3) / 4 is 0 or it stays as it is?

OpenStudy (anonymous):

Oh I think I get it... so [sqrt(3) / 4] * 2s?

OpenStudy (superdavesuper):

theres no real need to use the product rule....but if u insist on using it, then d/dx(k)=0 when k is a constant.

OpenStudy (superdavesuper):

that's right :)

OpenStudy (superdavesuper):

u can simplify it further between the /4 and *2 though...

OpenStudy (anonymous):

Ohhhh! I see now :) yes. I got s sqrt(3) /2 for final answer :)

OpenStudy (superdavesuper):

good job! :)

OpenStudy (anonymous):

there should be a variable in the answer since your s was squared

OpenStudy (anonymous):

Thanks for your help :D

OpenStudy (anonymous):

@Nameless s^2 turned 2s

OpenStudy (superdavesuper):

welcome

OpenStudy (anonymous):

ahh okay. then.. good job :)

OpenStudy (anonymous):

Thanks everyone :) :)

OpenStudy (anonymous):

[sqrt(3) / 4] * 2s - unless this was the original problem... there should be an s in your answer. d/ds(sqrt(3)/4)s^2 = sqrt(3)*s/2

OpenStudy (superdavesuper):

@nameless he did - " I got s sqrt(3) /2"

OpenStudy (anonymous):

oh.. i did not see the s out front. just trying to be thorough. :)

OpenStudy (superdavesuper):

w the Equation n draw tool down in OS, it has been driving us crazy for the last few days :(